OpenStudy (anonymous):

solve for x and y (x + 10)2 + y2 = 25

OpenStudy (anonymous):

is it your question \[(x+10)^2 +y^2=25\]

OpenStudy (anonymous):

@bekah2297 will u reply???

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thats my question.

OpenStudy (anonymous):

i need to know how to get the answer so i can do the other questions.

OpenStudy (anonymous):

is it correct???

OpenStudy (anonymous):

if yes then i will tell u how to get it

OpenStudy (anonymous):

im not sure,how did you get that.

OpenStudy (anonymous):

ok see \[(x+10)^2 + y^2 = 25 = 16 + 9 = 4^2 + 3^2\]

OpenStudy (anonymous):

it means that \[(x+10)^2 = 4^2\] and \[y^2=3^2\]

OpenStudy (anonymous):

are you getting it??????

OpenStudy (anonymous):

yes,thank you.

OpenStudy (anonymous):

ok on solving you will get x = -6 and y = 3 i m also not sure but i feel it is the correct way

OpenStudy (anonymous):

those arent any of my options.

OpenStudy (anonymous):

lol i too knew it may not be correct so asked u before anyways let me try once more can u tell me the options

OpenStudy (anonymous):

but i dont feel we can get answer to this question any other way because we have two variables and for that we require two equations